OpenStudy (anonymous):

what are ratios of 16:48

OpenStudy (anonymous):

Try dividing each side by 8

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

2:6

OpenStudy (anonymous):

but i need more ratios

OpenStudy (anonymous):

More the one? For that problem?

OpenStudy (anonymous):

yes thats what the problem is telling me

OpenStudy (anonymous):

Well what else do you think it could divide by

OpenStudy (anonymous):

um 2

OpenStudy (anonymous):

True you could do that..... But what about 4

OpenStudy (anonymous):

thats 4:12

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

now i need ratios for 36:48

OpenStudy (anonymous):

i dont like math ughhhhhhh

OpenStudy (anonymous):

lol math can be fun but also frustrating.

OpenStudy (anonymous):

true

OpenStudy (anonymous):

What is a common number that you can divide 36 by?

OpenStudy (anonymous):

Besides 2

OpenStudy (anonymous):

either 9 or 4

OpenStudy (anonymous):

and 3

OpenStudy (anonymous):

Ok now which one of those numbers can divide 48

OpenStudy (anonymous):

4, 12, 2, 6

OpenStudy (anonymous):

So all of those numbers can divide into both 36 and 48. Let's try 12.

OpenStudy (anonymous):

36/12=? 48/12=?

OpenStudy (anonymous):

thats 3:4

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

and now try 6

OpenStudy (anonymous):

36/6=? 48/6=?

OpenStudy (anonymous):

6:8

OpenStudy (anonymous):

yep

OpenStudy (texaschic101):

ratios can be written as fractions. 16:48 can be written as 16/48.....now just find equivalent fractions and you will have your answers

OpenStudy (anonymous):

um wats the greatest common factor of 10/12

OpenStudy (texaschic101):

lets see what numbers go evenly into 10.....1,2,5,10 now lets see what numbers go evenly into 12....1,2,3,4,6,12 now what numbers are common ? Which common number is the greatest ?

OpenStudy (anonymous):

2